### v1

@@ 3.4 Diffusion Model
Generative models have gained significant popularity in recent years, due to their ability to provide samples from complex distributions that characterize datasets ranging from images, audio, video, and language. The models must therefore learn the hidden structure of a distribution that represents the training dataset. In practice, all generative models follow two steps: first a sample is drawn from a distribution that is easy to sample from (e.g., a standard multivariate Gaussian) and second, the sample is mapped to a corresponding sample from the distribution of interest. Diffusion probabilistic models, or simply diffusion models (DM), which we use in this paper, sample from a standard Gaussian distribution and then in the second step learn a time reversed diffusion process. Reverse-time diffusion equations were first studied by Anderson [28] and then Haussmann and Pardoux [29]. These ideas, unkown to the original machine learning contributors, were introduced to machine learning for image generation by Sohl-Dickstein et al. [30]. Diffusion model capabilities in both conditional and unconditional image, audio, and video generation have since been greatly improved by Ho et al. [31], and Song and Ermon [32]. Following their success in image generation, diffusion models have been applied to areas like reinforcement learning [33, 34, 35], motion generation [36], and robotics [37].
Generation of samples using a diffusion model involves a time reversed diffusion process, but to train the model to learn this reverse process, a forward diffusion process is also required. Figure 5 illustrates the idea of the forward and reverse process. In particular, given samples $\boldsymbol{z}_{0}$ of the training dataset, these are evolved under a user prescribed forward diffusion, approximated with a fixed step integration scheme, to a final sample $\boldsymbol{z}_{N}$ after $N$ integration steps. The forward diffusion is chosen to yield final samples from a ”nice” distribution (i.e., one that will be easy to sample from during the generation process, such as a Gaussian distribution). The reverse process then learns the time-reversal of the forward diffusion equation, so that samples can be drawn from the ”nice” distribution and mapped to the desired and likely complex distribution.
In Fig. 5, this concept is illustrated for the global costate structure of a spacecraft trajectory optimization problem that will be described in the GTO Halo section of this paper. The trajectory corresponding to a selected data point from the costate distribution is visualized. In particular, the physical realization (i.e., simulation using pydylan) is illustrated at the bottom of the figure. As the sample is progressed through the reverse process, the feasibility of the trajectory is improved, until at $\boldsymbol{z}_{0}$ the sample has very little constraint error, such that an NLP solver can quickly converge to a feasible and then locally optimal solution. Additional details of these figures will be provided in the results section.
@@ 3.4.1 Forward Process
where $\mathbb{P}_{\alpha}$ is the distribution for the problem specified by the conditional parameter $\alpha$, $\boldsymbol{dw}_{t}$ is a standard Brownian motion, $\boldsymbol{\sigma}_{t}\succ 0,\forall t>0$, the dispersion coefficient (a positive definite matrix) and $\gamma>0$ the drift coefficient. Because $-\gamma$ provides dissipation and the Brownian motion is non-degenerate, the limiting distribution given by $\boldsymbol{z}_{t}$ is a standard Gaussian distribution. This stochastic differential equation is also a continuous-time Markov process.
@@ 3.4.2 Reverse Process
Assuming that the diffusion for any given step is sufficiently small (e.g., by choosing a sufficiently high number of integration steps and an appropriate dispersion coefficient schedule), it is appropriate to model the transition probability as a Gaussian distribution with conditional mean $\boldsymbol{\mu}_{\theta}$,
@@
and $\tilde{\beta}_{n}=(1-\overline{\alpha}_{n-1})/(1-\overline{\alpha}_{n})\cdot\beta_{n}$. The goal of training a diffusion model then becomes the standard problem of maximizing the likelihood of the data for the given probability distribution $p_{\theta}(\boldsymbol{z}_{0})$, which due to the Markovian property becomes a product of conditional Gaussian distributions. As is common in generative machine learning, a variational bound is instead extremized, since it has better numerical tractability. This variational bound is composed of two terms, a Kullback-Leibler (KL) divergence and a reconstruction loss. In the Gaussian setting, the KL divergence is computable with knowledge of the determinants of the covariances, which is most easily done when the covariances are only non-zero on the diagonal. If sample sizes are large enough, the KL term is less significant. Following this assumption, similar to Ho et al. [31] and Sohl-Dickstein et al. [30], we define the loss function to be extremized, and hence guide the learning of the DM, as just the reconstruction loss,
@@
where the expectation is over all possible diffusion steps in the reverse process, $\boldsymbol{\epsilon}_{n}$ is the known diffusion from the forward process, and $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{n},n)$ is the learned reverse diffusion noise. The entire step-by-step process is detailed in the work of Sohl-Dickstein et al. [30]. A neural network based on a U-Net architecture, originally introduced by Ronneberger et al. [39] is trained on this loss using a stochastic gradient descent algorithm.
@@ 4 Results
In this section, the strong performance of the diffusion model framework is presented for two example missions within the CR3BP system. The first mission involves a planar DRO transfer in the Jupiter-Europa system. This problem was chosen because its globally optimal solutions are well-characterized, having been identified as a smooth front of Pareto-optimal solutions by Russell [16]. To demonstrate the framework on a more complex scenario, we next examine a three-dimensional transfer in the Earth–Moon system. This second case is a modified version of the transfer solved by Li et al.[2], for which global optimality has not yet been confirmed for any solutions.
@@ 4.1.1 Problem Description
The first problem is a planar low-thrust transfer in the Jupiter-Europa CR3BP system. The spacecraft starts in an initial Distant Retrograde Orbit (DRO) at Europa and transfers to a lower DRO. The family of DROs around Europa evolves from planar, near circular orbits close to Europa to bigger, almost elliptical orbits with Europa at the center. Hence, this is a planar transfer, although the spacecraft is not explicitly restricted to stay in this plane by pydylan. The initial and final orbits, along with an example trajectory, are shown in Fig. 7.
@@
Figure 7: Linear relationship between the costates $\lambda_{r0,1}$ and $\lambda_{v0,2}$ for $\alpha=1.0$.
@@
Initial state $[\boldsymbol{r}_{0}^{T},\boldsymbol{v}_{0}^{T}]$ [NU] $[1.0752,0.0,0.0,0.0,-0.1499,0.0]$
Terminal state $[\boldsymbol{r}_{f}^{T},\boldsymbol{v}_{f}^{T}]$ [NU] $[1.0306,0.0,0.0,0.0,-0.0727,0.0]$
@@ 4.1.2 Data Generation
The goal for this problem is to train a diffusion model based on the conditional parameter $\alpha$, which describes the spacecraft’s maximum thrust level. The relationship between $T_{max}$ and $\alpha$ is defined by: $\alpha=T_{max}/(4.984\,\mathrm{N})$, with $\alpha\in[0.1,1.0]$. To train the model, a dataset is created, consisting of solutions for ten fixed thrust levels, which equidistantly cover the defined $\alpha$-range. The dataset includes 270,000 solutions, with 27,000 solutions per thrust level. Directly combining the preliminary screening algorithm with a small tolerance $\delta$ proves to be a highly efficient approach for quickly generating feasible solutions to this problem. Consequently, the strategy corresponding to the bottom path in Fig. 2 is used. Russell [16] demonstrated, that there exists a well-defined front of Pareto optimal solutions in the $\Delta v$-time-of-flight plane for this problem with $\alpha=1.0$. Solutions are labeled Pareto optimal if no other solutions have both a lower change in velocity ($\Delta v$) and a shorter time of flight. These solutions are specifically targeted in the data generation process, as they represent the most interesting trajectories from a mission design perspective. The global search parameters used for data generation are displayed in Table. 2. The ranges for $\varphi$, $\dot{\varphi}_{0}$, $\beta_{0}$, $\dot{\beta}_{0}$, $S_{0}$ and $\dot{S}_{0}$ were chosen based on the Pareto optimal solutions found by Russell [16] and further refined through test runs.
@@
[…]
@@
SNOPT mode optimal $\dot{S}_{0}$ [NU] $[-0.0022,0.004]$
The generated dataset primarily consists of solutions near the respective Pareto front for each $\alpha$, as demonstrated in Fig. 8.
@@
A subset of solutions for three different $\alpha$ values is displayed in the plane of the competing objectives, $\Delta v$ and shooting time $\tau_{s}$. The displayed Pareto front consist of the Pareto-approximate solutions from this subset for each thrust level. For $\alpha=1.0$, the Pareto front found by Russell [16] is reproduced. The lower thrust levels reveal a similar structure, with comparable $\Delta v$ and longer times of flight. For $\alpha=0.1$ the solutions tend to cluster more towards the upper part of the Pareto front, as generating solutions with longer flight times appears to be more challenging at very low thrust levels. However, these solutions are still valuable for mission design due to their comparatively shorter times of flight.
@@ 4.1.3 Solution Structure
Analyzing the solutions in the generated dataset reveals clusters in the costate space that vary depending on the conditional parameter $\alpha$. These clusters likely correspond to basins of locally optimal trajectories in the solution space.
@@
The new variable consolidates the linear dependency into a single dimension by capturing the maximum variance in those costates. The direction of the $\lambda^{\prime}$ axis is indicated by the black arrow in the bottom right plot of Fig. 7. The transformed costate variable permits a visualization of the two position and two velocity costates in a three-dimensional plot. The resulting curved hypersurface structure can be observed in Fig. 9 for three different thrust levels. The shape and location of these curved surfaces vary depending on the thrust level. A clear structure is evident for $\alpha=1.0$, but as the thrust level decreases, the structure becomes more diffuse, with the hypersurfaces barely recognizable for $\alpha=0.1$. Additionally, the shooting time $\tau_{s}$ is visualized through a color plot, revealing a trend where hypersurfaces with higher $\lambda^{\prime}$ values correspond to longer shooting times. However, many solutions with longer shooting times do not fit into the hypersurface structure.
@@ 4.1.4 Diffusion Model Training
The neural network used to model one step in the reverse process of the denoising diffusion probabilistic model from Ho et al. [31] is based on a U-Net architecture, originally introduced by Ronneberger et al. [39]. The diffusion model parameters that were used for training and sampling are presented in Table 3, with no claim for these parameters to be optimal for the given problem, since that would require an extensive hyperparameter study. For reproducibility and completeness we provide extra details on the network architecture and a description of all hyperparameters in the Appendix. In the following, a brief explanation of some of the trade-offs and comparisons made to justify the final parameter selection is given.
@@
Sequence length $L$ $6$ Channel number $C$ $1$
@@
Encoder layer dimensions $[256,256,512]$ Timesteps $N$ 1,000
@@
The batch size represents the subset of training data used during each step of the training process. It controls how much GPU computational capacity and memory are used, directly impacting the training time. Increasing the batch size reduces training time up to the point where GPU memory usage reaches its limit. However, a larger batch size also reduces variance in the generated samples, leading to lower sample diversity. The size of the U-Net influences the network’s capacity to capture and represent the complexity within the data structure. It is mainly determined through the number and dimensions of the encoder layers and embedded class layers, as well as the dimension of the time embedding. Increasing the number of layers or their dimensions also leads to longer training and sampling times.
The diffusion model is implemented and trained in PyTorch [41]. Before training begins, the data is randomly split into a training set and a validation set in a 9:1 ratio. Training runs for a fixed number of epochs, where each epoch involves iterating through the entire training dataset. The number of epochs is selected based on the validation loss, as shown in Fig. 11. The moving average does not significantly decrease after 150 epochs, indicating that additional training beyond 200 epochs is unnecessary. Choosing the number of diffusion timesteps, $N$, involves a trade-off between sampling quality and sampling time. The chosen value of $N$ = 1,000 results in a sampling time of $330.0$ s for 50,000 samples on a NVIDIA A100-SXM4-40GB GPU. The diversity of the generated samples is controlled by the guidance strength $w$. If this value is set too high, the model primarily targets the largest clusters in the solution space and fails to generate initializations in the less densely populated regions. Conversely, setting $w$ too low reduces the model’s ability to incorporate classifier information during sample generation.
@@
Figure 11: Hypersurface structure in the costate space of the training data in comparison to diffusion model samples.
Figure 12: Structure in the costate space after different diffusion timesteps for thrust level $\alpha=1.0$. Each plot displays a subset of 5,000 solutions. Note the different axis and colorbar scaling in the first plot, which shows the Gaussian noise at $n=1$,000.
The hyperparemeters of the model were chosen based on a visual evaluation of the model’s ability to reproduce the structure in the costate space for $\alpha=1.0$. Since $\alpha=1.0$ is part of the training dataset, a well-chosen set of parameters should enable the model to accurately generate a similar structure, starting from Gaussian noise. This process is visualized for the final choice of parameters in Fig. 12. The top left plot displays the initial data points sampled from a multivariate standard normal distribution. The data is centered in the middle of the bounds used for normalization, but extends well past those bounds, resulting in larger costate values and unreasonable shooting times $\tau_{s}$. Over the course of $N$ = 1,000 timesteps, the structure in the data is gradually reconstructed from the initial noise. The dataset corresponding to the prediction of $\boldsymbol{z}_{0}$ after $n$ timesteps is shown for selected values of $n$. For this visualization, $\boldsymbol{z}_{0}$ was calculated by plugging $\boldsymbol{z}_{n}$ and the predicted noise from the neural network into Eq. (24). The displayed process here is equivalent to the visualization in Fig. 5. A comparison of diffusion model samples with the training data structure in Fig. 11 demonstrates that the chosen parameters produce valid results. A more detailed explanation of the hyperparameter choices, along with results from models tested with different parameter settings for this problem, is provided in the Master’s thesis of Gräbner[42].

### v2

@@ 3.4 Diffusion Model
Generative models have gained significant popularity in recent years, due to their ability to provide samples from complex distributions that characterize datasets ranging from images, audio, video, and language. The models must therefore learn the hidden structure of a distribution that represents the training dataset. In practice, all generative models follow two steps: first a sample is drawn from a distribution that is easy to sample from (e.g., a standard multivariate Gaussian) and second, the sample is mapped to a corresponding sample from the distribution of interest. Diffusion probabilistic models, or simply diffusion models (DM), which we use in this paper, sample from a standard Gaussian distribution and then in the second step learn a time reversed diffusion process. Reverse-time diffusion equations were first studied by Anderson [31] and then Haussmann and Pardoux [32]. These ideas, unknown to the original machine learning contributors, were introduced to machine learning for image generation by Sohl-Dickstein et al. [33]. Diffusion model capabilities in both conditional and unconditional image, audio, and video generation have since been greatly improved by Ho et al. [34], and Song and Ermon [35]. Following their success in image generation, diffusion models have been applied to areas like reinforcement learning [36, 37, 38], motion generation [39], and robotics [40].
Generation of samples using a diffusion model involves a time-reversed diffusion process, but to train the model to learn this reverse process, a forward diffusion process is also required. Figure 5 illustrates the idea of the forward and reverse process. In particular, given samples $\boldsymbol{z}_{0}$ of the training dataset, these are evolved under a user-prescribed forward diffusion, approximated with a fixed step integration scheme, to a final sample $\boldsymbol{z}_{N}$ after $N$ integration steps. The forward diffusion is chosen to yield final samples from a ”nice” distribution (i.e., one that will be easy to sample from during the generation process, such as a Gaussian distribution). The reverse process then learns the time-reversal of the forward diffusion equation, so that samples can be drawn from the ”nice” distribution and mapped to the desired and likely complex distribution.
@@
In Fig. 5, this concept is illustrated for the global costate structure of a spacecraft trajectory optimization problem that will be described in the GTO Halo section of this paper. The trajectory corresponding to a selected data point from the costate distribution is visualized. In particular, the physical realization (i.e., simulation using pydylan) is illustrated at the bottom of the figure. As the sample is progressed through the reverse process, the feasibility of the trajectory is improved, until at $\boldsymbol{z}_{0}$ the sample has very little constraint error, such that an NLP solver can quickly converge to a locally optimal solution. Additional details of these figures will be provided in the results section.
@@ 3.4.1 Forward Process
where $\mathbb{P}_{\alpha}$ is the distribution for the problem specified by the conditional parameter $\alpha$, $\boldsymbol{dw}_{t}$ is standard Brownian motion, $\boldsymbol{\sigma}_{t}\succ 0,\forall t>0$, the dispersion coefficient (a positive definite matrix) and $\gamma>0$ the drift coefficient. Because $\gamma$ provides dissipation and the Brownian motion is non-degenerate, the limiting distribution given by $\boldsymbol{z}_{t}$ is a standard Gaussian distribution. This stochastic differential equation is also a continuous-time Markov process.
@@ 3.4.2 Reverse Process
A neural network is a flexible mathematical model built from layers of simple processing units (“neurons”). The parameters of the network, which we denote by $\theta$ in Eq. (25), are tuned during training to specify their connection weights and learn to transform inputs into accurate outputs. Assuming that the diffusion for any given step is sufficiently small (e.g., by choosing a sufficiently high number of integration steps and an appropriate dispersion coefficient schedule), it is appropriate to model the transition probability as a Gaussian distribution with conditional mean $\boldsymbol{\mu}_{\theta}$,
@@
and $\tilde{\beta}_{n}=(1-\overline{\alpha}_{n-1})/(1-\overline{\alpha}_{n})\cdot\beta_{n}$. The neural network takes $\boldsymbol{z}_{n}$ and $n$ as input and outputs $\boldsymbol{\mu}_{\theta}$. The goal of training a diffusion model then becomes the standard problem of maximizing the likelihood of the data for the given probability distribution $p_{\theta}(\boldsymbol{z}_{0})$, which due to the Markovian property becomes a product of conditional Gaussian distributions. As is common in generative machine learning, a variational bound is instead extremized, since it has better numerical tractability. This variational bound is composed of two terms, a Kullback-Leibler (KL) divergence and a reconstruction loss. In the Gaussian setting, the KL divergence is computable with knowledge of the determinants of the covariances, which is most easily done when the covariances are only non-zero on the diagonal. If sample sizes are large enough, the KL term is less significant. Following this assumption, similar to Ho et al. [34] and Sohl-Dickstein et al. [33], we define the loss function to be extremized, and hence guide the learning of the DM, as just the reconstruction loss,
@@
where the expectation is over all possible diffusion steps in the reverse process, $\boldsymbol{\epsilon}_{n}$ is the known diffusion from the forward process, and $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{n},n)$ is the learned reverse diffusion noise. The entire step-by-step process is detailed in the work of Sohl-Dickstein et al. [33]. A neural network based on a U-Net architecture, originally introduced by Ronneberger et al. [42] is trained on this loss using a stochastic gradient descent (SGD) algorithm. In SGD, the model’s parameters are updated iteratively using an approximation to the true expected gradient by computing an averaged one using a small, randomly selected mini-batch of gradients associated with the training data. In each iteration, the mini-batch is fed forward through the network to produce predictions, the loss between these predictions and the ground truth is calculated, and the resulting error gradients are back-propagated to adjust the network weights.
@@ 4 Results
In this section, the strong performance of the diffusion model framework is presented for two example missions within the CR3BP system. The first mission involves a planar DRO transfer in the Jupiter-Europa system. This problem was chosen because it has a smooth front of Pareto-optimal solutions found by Russell [19], which are likely to be good numerical approximations of the true Pareto front. To demonstrate the framework on a more complex scenario, we next examine a three-dimensional transfer in the Earth–Moon system. This second case is a modified version of the transfer solved by Li et al.[2], for which no clear Pareto front has been found yet.
@@ 4.1.1 Problem Description
The first problem is a planar low-thrust transfer in the Jupiter-Europa CR3BP system. It derives from the Jupiter Icy Moons Orbiter (JIMO) mission concept, which was canceled in 2005. The spacecraft starts in an initial Distant Retrograde Orbit (DRO) at Europa and transfers to a lower DRO. The family of DROs around Europa evolves from planar, near circular orbits close to Europa to bigger, almost elliptical orbits with Europa at the center. Hence, this is a planar transfer, although the spacecraft is not explicitly restricted to stay in this plane by pydylan. The initial and final orbits, along with an example trajectory, are shown in Fig. 7.
@@
Figure 7: Linear relationship between the costates $\lambda_{r0,1}$ and $\lambda_{v0,2}$ for $\alpha=1.0$ $(T_{max}=4.984\,\mathrm{N})$.
@@
Initial state $\boldsymbol{r}_{0}^{T},\boldsymbol{v}_{0}^{T}] [$NU] [1.0752,0.0,0.0,0.0,0.1499,0.0][1.0752,0.0,0.0,0.0,-0.1499,0.0]
Terminal state $\boldsymbol{r}_{f}^{T},\boldsymbol{v}_{f}^{T}] [$NU] [1.0306,0.0,0.0,0.0,0.0727,0.0][1.0306,0.0,0.0,0.0,-0.0727,0.0]
@@ 4.1.2 Data Generation
The goal for this problem is to train a diffusion model based on the conditional parameter $\alpha$, which describes the spacecraft’s maximum thrust level. The relationship between $T_{max}$ and $\alpha$ is defined by: $\alpha=T_{max}/(4.984\,\mathrm{N})$, with $\alpha\in[0.1,1.0]$. To train the model, a dataset is created, consisting of solutions for ten fixed thrust levels, which equidistantly cover the defined $\alpha$-range. The dataset includes 270,000 solutions, with 27,000 solutions per thrust level. Directly combining the preliminary screening algorithm with a tolerance $\delta=1\times 10^{-4}$ [NU] proves to be a highly efficient approach for quickly generating feasible solutions to this problem. Consequently, the strategy corresponding to the bottom path in Fig. 2 is used. Although this tolerance is coarse by the standards of mission design, it allows us to capture the global solution distribution efficiently. Russell [19] demonstrated that there exists a well-defined front of Pareto optimal solutions in the $\Delta v$-time-of-flight plane for this problem with $\alpha=1.0$. Solutions are labeled Pareto optimal if no other solutions have both a lower change in velocity ($\Delta v$) and a shorter time of flight. These solutions are specifically targeted in the data generation process, as they represent the most interesting trajectories from a mission design perspective. The global search parameters used for data generation are displayed in Table. 2. The ranges for $\varphi$, $\dot{\varphi}_{0}$, $\beta_{0}$, $\dot{\beta}_{0}$, $S_{0}$ and $\dot{S}_{0}$ were chosen based on the Pareto optimal solutions found by Russell [19] and further refined through test runs.
@@
Global search tolerance $\delta$ [NU] $1\text{\times}{10}^{-4}$ $\varphi_{0}$ [rad] $\pi-0.012,\pi+0.01]$
Feasibility tolerance [NU] $1\text{\times}{10}^{-4}$ $\dot{\varphi}_{0}$ [rad/TU] [0.02,0.025][-0.02,0.025]
@@
SNOPT mode optimal $\dot{S}_{0}$ [NU] [0.0022,0.004][-0.0022,0.004]
@@
The generated dataset primarily consists of solutions near the respective Pareto front for each $\alpha$, as demonstrated in Fig. 8. A subset of solutions for three different $\alpha$ values is displayed in the plane of the competing objectives, $\Delta v$ and shooting time $\tau_{s}$. The displayed Pareto front consists of the Pareto-approximate solutions from this subset for each thrust level. For $\alpha=1.0$, the Pareto front found by Russell [19] is reproduced. The lower thrust levels reveal a similar structure, with comparable $\Delta v$ and longer times of flight. For $\alpha=0.1$ the solutions tend to cluster more towards the upper part of the Pareto front, as generating solutions with longer flight times appears to be more challenging at very low thrust levels. However, these solutions are still valuable for mission design due to their comparatively shorter times of flight.
@@ 4.1.3 Solution Structure
Analyzing the solutions in the generated dataset reveals clusters in the costate space that vary depending on the thrust magnitude $T_{max}$. These clusters likely correspond to basins of locally optimal trajectories in the solution space.
@@
The new variable consolidates the linear dependency into a single dimension by capturing the maximum variance in those costates. The direction of the $\lambda^{\prime}$ axis is indicated by the black arrow in the bottom right plot of Fig. 7. The transformed costate variable permits a visualization of the two position and two velocity costates in a three-dimensional plot.
The resulting curved hypersurface structure can be observed in Fig. 9 for three different thrust levels. The shape and location of these curved surfaces vary depending on the thrust level. A clear structure is evident for $\alpha=1.0$, but as the thrust level decreases, the structure becomes more diffuse, with the hypersurfaces barely recognizable for $\alpha=0.1$. Additionally, the shooting time $\tau_{s}$ is visualized through a color plot, revealing a trend where hypersurfaces with higher $\lambda^{\prime}$ values correspond to longer shooting times. However, many solutions with longer shooting times do not fit into the hypersurface structure.
@@ 4.1.4 Diffusion Model Training
The neural network used to model one step in the reverse process of the denoising diffusion probabilistic model from Ho et al. [34] is based on a U-Net architecture, originally introduced by Ronneberger et al. [42]. A U-Net consists of an encoder and a decoder. The encoder compresses the input data into a lower-dimensional hidden representation, resulting in a bottleneck layer. After the bottleneck, the decoder restores this representation to the original input dimension, with residual connections from the encoder to the decoder enhancing gradient flow. The diffusion model parameters that were used for training and sampling are presented in Table 3, with no claim for these parameters to be optimal for the given problem, since that would require an extensive hyperparameter study. For reproducibility and completeness we provide extra details on the network architecture and a description of all hyperparameters in the Appendix. In the following, a brief explanation of some of the trade-offs and comparisons made to justify the final parameter selection is given.
@@
Sequence length $L$ 66 Channel number $C$ 11
@@
Encoder layer dimensions $[256,256,512]$ Timesteps NN 1,000
@@
The batch size represents the subset of training data used during each step of the training process. It controls how much GPU computational capacity and memory are used, directly impacting the training time. Increasing the batch size reduces training time until the GPU’s memory limit is reached. However, a larger batch size also reduces variance in the generated samples, leading to lower sample diversity. The size of the U-Net influences the network’s capacity to capture and represent the complexity within the data structure. It is mainly determined through the number and dimensions of the encoder layers and embedded class layers, as well as the dimension of the time embedding. Increasing the number of layers or their dimensions also leads to longer training and sampling times.
The diffusion model is implemented and trained in PyTorch [44]. Before training begins, the data is randomly split into a training set and a validation set in a 9:1 ratio. The training set is the portion on which the model actually learns and updates its weights, while the validation set checks its performance on unseen data. Training runs for a fixed number of epochs, where each epoch involves iterating through the entire training dataset. The number of epochs is selected based on the validation loss, as shown in Fig. 11. The moving average does not significantly decrease after 150 epochs, indicating that additional training beyond 200 epochs is unnecessary. Choosing the number of diffusion timesteps, $N$, involves a trade-off between sampling quality and sampling time. The chosen value of $N$ = 1,000 results in a sampling time of $330.0$ s for 50,000 samples on a NVIDIA A100-SXM4-40GB GPU. The diversity of the generated samples is controlled by the guidance strength $w$. If this value is set too high, the model primarily targets the largest clusters in the solution space and fails to generate initializations in the less densely populated regions. Conversely, setting $w$ too low reduces the model’s ability to incorporate classifier information during sample generation.
@@
Figure 11: Hypersurface structure in the costate space of the training data in comparison to diffusion model samples for  $\alpha=1.0\,(T_{max}=4.98\,\mathrm{N})$.
Figure 12: Structure in the costate space after different diffusion timesteps for thrust level $\alpha=1.0\,(T_{max}=4.98\,\mathrm{N})$. Each plot displays a subset of 5,000 solutions. Note the different axis and colorbar scaling in the first plot, which shows the Gaussian noise at n=1n=1,000.
The hyperparameters of the model were chosen based on a visual evaluation of the model’s ability to reproduce the structure in the costate space for $\alpha=1.0$. While this is no quantifiable comparison, it is the best option to ensure that the costate structure is fully captured by our model. A comparison of how this structure is captured by models with varying parameters is presented in the appendix. The reference model is shown in the first panel and each remaining panel varies in exactly one parameter, highlighting how that parameter impacts the distribution learned by the model. Since $\alpha=1.0$ is part of the training dataset, a well-chosen set of parameters should enable the model to accurately generate a similar structure, starting from Gaussian noise. This process is visualized for the final choice of parameters in Fig. 12. The top left plot displays the initial data points sampled from a multivariate standard normal distribution. The data is centered in the middle of the bounds used for normalization, but extends well past those bounds, resulting in larger costate values and unreasonable shooting times $\tau_{s}$. Over the course of $N$ = 1,000 timesteps, the structure in the data is gradually reconstructed from the initial noise. The dataset corresponding to the prediction of $\boldsymbol{z}_{0}$ after $n$ timesteps is shown for selected values of $n$. For this visualization, $\boldsymbol{z}_{0}$ was calculated by plugging $\boldsymbol{z}_{n}$ and the predicted noise from the neural network into Eq. (24). The displayed process here is equivalent to the visualization in Fig. 5. A comparison of diffusion model samples with the training data structure in Fig. 11 demonstrates that the chosen parameters produce valid results. Slight modifications of the hyperparameters chosen in this paper may yield slightly improved results in the following sections, though are not expected to change results substantially. A comprehensive hyperparameter study is beyond the scope of this work. A more detailed explanation of the hyperparameter choices, along with results from models tested with different parameter settings for this problem, is provided in the Master’s thesis of Gräbner[45].